OpenStudy (anonymous):

foci(17+/,0), vertices (+/8,0) c^2 = a^2 – b^2 c = 17 and a = 8 c^2= 289 and a^2 = 64 a^2= c^2 – b^2 64 = 289 – 225 Equation of ellipse x^2/225+y^2/289=1 Is this right?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@texaschic101

OpenStudy (phi):

shouldn't the vertices be outside of the focus points ? In other words, are you sure you didn't switch them ?

OpenStudy (anonymous):

Probably. If it makes a difference, then what should it be?

OpenStudy (phi):

either that, or you have a hyperbola ? What information are you given?

OpenStudy (anonymous):

It is, Write an equation of a hyperbola with the given foci and vertices.

OpenStudy (phi):

then it's clear it's not an ellipse. so your answer has to be tweaked.

OpenStudy (phi):

This gives the details http://www.purplemath.com/modules/hyperbola.htm

OpenStudy (anonymous):

Am I suppose to use a different equation?

OpenStudy (phi):

Based on the picture at the posted link, and the equation they give, the equation should be \[ \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1\] we need to find a and b a is the distance from the center to the vertex. the center is exactly between the 2 vertexes, i.e. between (-8,0) and (+8,0) , which is (0,0)

OpenStudy (phi):

so you did this part correctly c = 17 and a = 8 c^2= 289 and a^2 = 64 to find b (for a hyperbola) the link says use this formula \[ c^2 = a^2 + b^2 \] or , solving for b^2, \[ b^2 = c^2 - a^2\]

OpenStudy (phi):

can you find b ?

OpenStudy (anonymous):

50625?

OpenStudy (phi):

I think that is b^2

OpenStudy (anonymous):

then b is 225

OpenStudy (phi):

no, b^4

OpenStudy (anonymous):

b^4?

OpenStudy (anonymous):

Am I looking for b^4 or am I writing the equation?

OpenStudy (phi):

You should be finding b using b^2 = c^2 - a^2 c=17, a=8

OpenStudy (phi):

or just find b^2 (don't bother taking the square root to find b)

OpenStudy (anonymous):

I did find b2, you said that's wrong b=225 b^2=50625, otherwise I have no idea what you're referring to.

OpenStudy (phi):

b^2 is 225 (and b is 15).

OpenStudy (anonymous):

Oh. . .I don't know where 15 came from.

OpenStudy (anonymous):

Oh nevermind sqrt225.

OpenStudy (phi):

How did you find b^2. what steps did you do?

OpenStudy (anonymous):

289-64. . .>~>

OpenStudy (phi):

and that gives 225 in other words b^2 is 225 I hope you are not getting confused about what is squared and what isn't. But anyways, a = 8 and a^2 is 64 b is 15, and b^2 is 225 use those numbers in \[ \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1 \]

OpenStudy (anonymous):

You confused me by asking me what b^2 was, and I already told you. xD

OpenStudy (anonymous):

x^2/15 - y^2/225=1?

OpenStudy (phi):

is a^2 = 15 ? because you want a^2 under the x^2

OpenStudy (anonymous):

The other way around?

OpenStudy (phi):

no. First, look at what you found so far: what is a? a is 8 what is a^2 =a*a ? it is 64 what is b? b is 15. b^2 = b*b = 225 Now look at the equation \[ \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1 \] replace a^2 with the number for a^2 replace the b^2 with its number

OpenStudy (anonymous):

Ohhh, kekeke. x^2/64-y^2/225

OpenStudy (phi):

almost. equations have = signs in them

OpenStudy (anonymous):

=1

OpenStudy (phi):

Here is your hyperbola

OpenStudy (anonymous):

. . .Eh, that sucks. Should have used the calculator. Thanks as always.

OpenStudy (phi):

Working with the numbers "by hand" is the only way to get reasonably good at it. And though it might not seem so, using numbers is a good thing to be able to do.

OpenStudy (anonymous):

Yup, if phi says so. ^^